OpenStudy (anonymous):

we have 3 balls of color x and 4 balls of color y in an urn, what is the probability of drawing 2 balls of same color without replacement??

OpenStudy (anonymous):

i have it forgotten :(

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the probability that both balls are color x

OpenStudy (anonymous):

C(2,3)/ C(2,7) ???

jimthompson5910 (jim_thompson5910):

you have the notation mixed up

jimthompson5910 (jim_thompson5910):

it should be C(3,2)/C(7,2)

jimthompson5910 (jim_thompson5910):

what does that evaluate to

OpenStudy (anonymous):

1/7

jimthompson5910 (jim_thompson5910):

so the probability of picking 2 balls that are color x is 1/7 what is the probability of picking 2 balls that are color y?

ganeshie8 (ganeshie8):

Is there some trick.. isnt it just \(\frac{3}{7} \frac{2}{6} + \frac{4}{7}\frac{3}{6}\)

OpenStudy (anonymous):

y: 2/7

jimthompson5910 (jim_thompson5910):

so the probability of picking 2 balls that are color y is 2/7 add them up now 1/7+2/7 = 3/7 this means the probability of selecting 2 balls of the same color is 3/7

OpenStudy (campbell_st):

an easy way to do this is a tree diagram  so you can have P(x + x) + P( y + y) = 3/7*2/6 + 4/7*3/6